OpenStudy (aaronandyson):

Circle A has 1/3 the radius of circle B, and circle A rolls one trip around circle B. How many times will circle A revolve in total?

OpenStudy (sshayer):

circumference=2 pi r it depends on r only. \[\frac{ circumference ~of~B }{ circumference ~of~A }=\frac{ 2 \pi r }{ 2 \pi \left( \frac{ 1 }{ 3 }r \right) }\] =?

OpenStudy (mayankdevnani):

but 3 is incorrect

OpenStudy (sshayer):

what is the answer or options?

OpenStudy (mayankdevnani):

i dun know the answer :( he said 3 is incorrect

OpenStudy (aaronandyson):

(a) 3/2 (b) 3 (c) 6 (d) 9/2 (e) 9

OpenStudy (aaronandyson):

Remember its REVOLVING.

OpenStudy (robtobey2):

(b)

OpenStudy (mayankdevnani):

oh! it's revolving isn't 4 correct ?

OpenStudy (aaronandyson):

YES!

OpenStudy (mayankdevnani):

but your choices ?

OpenStudy (aaronandyson):

The guy who prepared the question FORGOT to include 4 as the correct answer.

OpenStudy (mayankdevnani):

ohk! btw i like it

OpenStudy (sshayer):

i think you solved like this.  \[\frac{ 2 \pi (\frac{ 4 }{ 3 })r }{ 2 \pi \left( \frac{ 1 }{ 3 } r\right) }\]